OpenStudy (anonymous):

Which of the following points lie in the solution set to the following system of inequalities? y < -3x + 3 y < x + 2

OpenStudy (anonymous):

To get the points we should make two equations equal; here: -3x+3=x+2 -3x-x=2-3 -4x=-1 x=1/4.